OpenStudy (anonymous):

whats the focus of the parabola y=1/8(x-2)^2 -4

OpenStudy (anonymous):

can anyone help me!?

OpenStudy (amistre64):

it we put it into the geometric form of: (y-k)^2 = 4a(x-h) itll be easier to deal with

OpenStudy (anonymous):

how do you do that?

OpenStudy (amistre64):

by moving things about :)

OpenStudy (anonymous):

but what are all those variables?

OpenStudy (amistre64):

they define the center, and the distance from the vertex to the focus/directix

OpenStudy (amistre64):

y=1/8(x-2)^2 -4 +4 +4 y+4 =1/8(x-2)^2 *8 *8 8(y+4) = (x-2)^2 (4*2) (y+4) = (x-2)^2

OpenStudy (amistre64):

trying to remember if i sholda moved the 8 :/

OpenStudy (anonymous):

ohh okay! so then whats the focus?

OpenStudy (amistre64):

(y+4) =(1/8) (x-2)^2 (y+4) = 4 (1/32) (x-2)^2 that might be better

OpenStudy (amistre64):

the center and vertex of a parabola are the same point; (2,-4) if we read it form the equation

OpenStudy (amistre64):

the focus is "a" away from the center in the direction of the open part; in this case add 1/32 to the center y

OpenStudy (amistre64):

and to make sure im remembering it right, well chk with the wolf

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D1%2F8%28x-2%29%5E2+-4 pfft, i was right the first time with a=2 .....

OpenStudy (anonymous):

So that would make the focus (2, -2) right?

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

yay! thankyou!

OpenStudy (amistre64):

to recap the correct rendition :) y=1/8(x-2)^2 -4 +4 +4 y+4 =1/8(x-2)^2 *8 *8 8(y+4) = (x-2)^2 4(2) (y+4) = (x-2)^2 4 a (y -k) = (x-h)^2 center is (h,k), and we add "a" good luck